5. Determine the BSA, express your answer to the nearest hundredth. For a child weighing 92 pounds 12 ounces and who measures 35 inches.

media%2Fc28%2Fc285f5b0-9d32-4bf9-94fd-55
Answer #1

1 One dose =100mg/m^2/dose*2=200mg/m^2

2 Convert weight in pounds=2.2lbs/1kg=92lbs/x=92/2.2x=X=41.8kg

Convert height in inches to cm= Multiply by 2.54=35 inch*2.54=88.9cm

BSA*DOSAGE=200*1.91=382mg/m^2

3 yes it needs to be questioned as the single dosage is 200mg/m^2

4Yes it needs to be questioned as it exceeds the dosage of 382mg/m^2